OpenStudy (anonymous):

what value of y makes y/4=1.3/2 a propotion

OpenStudy (anonymous):

y/4=1.3/2a Multiplying by 4 both sides we get y/4x4=1.3a/2ax4 y=5.2/2a y=2.6/a